How to solve this question that what is the value of 1.008^(1/8)?

Grade:12th pass

2 Answers

Deepak Kumar Shringi
5 years ago
to find this value you should use log tables
Aabha Misra
5 years ago
Use binomial theorem,
 
General expression is : (1 + x)n = 1 + nx
 
The expression will become (1 + 1/8 * 0.008) = 1.001